OpenStudy (anonymous):

Pierre hit a baseball, and the flight of the ball was modeled by the following equation: h = -16t2 + 61t + 3. After how many seconds will the ball fall to the ground, if it is not caught? Round your answer to the nearest tenth. (a) 4.0 seconds (b) 3.8 seconds (c) 3.7 seconds (d) 3.9 seconds

OpenStudy (campbell_st):

you need to find when h = 0 ( the ball is on the ground) then \[0 = -16t^2 + 61t + 3\] use the general quadratic formula to find t \[t = \frac{-61 \pm \sqrt{61^2 - 4\times -16\times3}}{2\times-16}\] only consider the positive t value in your solution.

OpenStudy (anonymous):

When the height is zero...